Find the final amount of money in an account if $4,700 is deposited at 3% interest compounded quarterly (every 3 months) and the money is left for 6 years.
The final amount is $_____. Round answer to 2 decimal places

Answer by Theo(13342)

f = p * (1+r)^n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

there are 4 quarters in a year.

3% interest per year compounded quarterly gives you 3% / 4 = .75% interest rate per quarter.

.75% interest rate per quarter is equal to .0075 interest rate per quarter.

.0075 is the decimal equivalent of .75%.

the decimal equivalent of the percent is what is used in the formula.

6 years * 4 quarters per year = 24 quarters.

you are given that p = 4700

the equation of f = p * (1+r)^n becomes:

f = 4700 * (1.0075)^24

solve for f to get:

f = 5623.143588

if you are given the yearly interest rate and the number of years, then the formula can also be shown as:

f = p * (1 + r/c) ^ (nc)

f is the future value
p is the present value
r/c is the yearly interest rate divided by the number of compounding periods per year.
nc is the number of years multiplied by the compounding periods per year.

your formula would have become f = 4700 * (1 + .03/4) ^ (6*4)

this would have resulted in f = 5623.143588 which is the same answer as what was previously calculated above.